# Weyl’s Law for compact rank one symmetric spaces

### v1

@@ 1. Introduction
$N_{M}(\lambda)$ :=#$:=\#\{\textnormal{eigenvalues of $\Delta_{M}$}\leq\lambda^{2}\}.$
@@ 3. Setup and notation
$\Lambda_{M}$ :=$:=\{\sqrt{\alpha_{M}(k)}:k=0,1,2,\ldots\},\qquad\textnormal{where}\qquad\alpha_{M}(k):=A_{M}k^{2}+B_{M}k+C_{M}.$ (5)
@@
Assume that the multiplicity of these eigenvalues are given by a degree $d-1$ polynomial $\mathsf{R}_{M}(k)$. Define
@@
where $\varphi(t)$ is a polynomial of degree at most $d-3$ and $C$ is some constant.
@@ 4. A lattice counting result by Iosevich and Wyman
The following result is an immediate consequence of the previous theorem.
@@
Let $d_{1},d_{2},\ldots,d_{n}\in\mathbb{Z}^{\geq 2}$ and $n\geq 2$. Let $y_{1},\ldots,y_{n}$ be fixed $n$ real numbers and $P_{1},P_{2},\ldots P_{n}$ be fixed $n$ positive real numbers. Then
@@
Here, we use the norm $p$ defined on $\mathbb{R}^{n}$ by
@@
The crucial fact is that the ellipsoid $p^{-1}(1)$ has non vanishing Gaussian curvature. Let us now use the above result to prove theorem 1.2 of this paper.
@@ 5. Error can be polynomially improved for products of $\mathsf{W}$-manifolds: proof
Note that the sum total of the powers of all the $m_{i}$ is $d-n$.
@@
where the sum of the total power $e_{1}+\ldots e_{n}$ is at most $d-n-2$ (due to the condition 10). In particular, the sum total of the powers of $m_{i}$ in the summand of each of the terms of $T_{2}$ is at least $2$ less than the sum total of the powers of the $m_{i}$ in the summand of $T_{1}$.
@@
Similarly, we again apply theorem 4.2 and the binomial theorem to each of the terms of $T_{2}$. Doing that, we conclude that $T_{2}$ will be given by the right hand side of (17) with $d$ replaced by $d-2$, i.e.
@@ 6.1. Odd dimensional spheres
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree $22d-$ 2. Now,
@@ 6.2. Even dimensional spheres
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree $12d-$ 1. Now,
@@ 6.3. Odd dimensional real projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree $22d-$ 2. Now,
@@ 6.4. Even dimensional real projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree $12d-$ 1. Now,
@@ 6.5. Complex projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree $12d-$ 1. Now,
@@ 6.6. Quaternionic projective spaces
Let us now move on to the quaternionic projective spaces $M=\mathbb{HP}^{d-1}$, with $d\geq 2$. We are referring to the (quaternionic) dimension as $d-1$ as opposed to $d$ in order to make it easy to follow the notation of [4]. Notice that the real dimension of $\mathbb{HP}^{d-1}$ is $44d-$ 4. Following [4, page 10, theorem 5.2], we conclude that
@@
Notice that $\mathsf{R}_{M}(k)$ is a polynomial of degree $54d-$ 5. Now,
@@ 7. Error is sharp for a single $\mathsf{W}$-manifold: proof
where $\mathcal{R}_{k-2}(n)$ is a polynomial in $n$ of degree at most $k-2$.
@@
$1\displaystyle 2^{k}-1^{k}=k\cdot 1^{k-1}+\frac{k(k-1)}{2}\cdot 1^{k-2}+\cdots+$ 1 .
@@
Corollary 7.2.
Let us define the following function $N:\mathbb{R}\longrightarrow\mathbb{R}$, given by
$N(\lambda)$ $:=\sum_{i=1}^{[\lambda]}i^{k-1}$
$=1^{k-1}+2^{k-1}+\ldots+[\lambda]^{k-1},$ (56)
where $k\in\mathbb{Z}^{\geq 2}$. Then,
$N(\lambda)$ $=\frac{1}{k}[\lambda]^{k}+\frac{1}{2}[\lambda]^{k-1}+\mathcal{R}_{k-2}([\lambda]),$
where $\mathcal{R}_{k-2}([\lambda])$ is a polynomial in $[\lambda]$ of degree at most $k-2$.
Proof.
Follows immediately from proposition 7.1. ∎
Let us now make a couple of definitions, which will make it easier to state the remaining results.
Definition 7.3.
Let $N,M:\mathbb{R}\rightarrow\mathbb{R}$. We will say that $N$ is asymptotically equivalent to $M$ to first order (which will be written as $N\sim_{1}M$) if there exists $d\in\mathbb{Z}^{\geq 1}$ such that
$\lim_{\lambda\rightarrow\infty}\frac{N(\lambda)}{\lambda^{d}}=c_{1}=\lim_{\lambda\rightarrow\infty}\frac{M(\lambda)}{\lambda^{d}},$
where $c_{1}$ is a non-zero real number.
Definition 7.4.
Let $N,M:\mathbb{R}\rightarrow\mathbb{R}$. We will say that $N$ is asymptotically equivalent to $M$ to second order (which will be written as $N\sim_{2}M$) if there exists $d\in\mathbb{Z}^{\geq 1}$ such that
$\lim_{\lambda\rightarrow\infty}\frac{N(\lambda)}{\lambda^{d}}=c_{1}=\lim_{\lambda\rightarrow\infty}\frac{M(\lambda)}{\lambda^{d}},$
and
$\lim_{\lambda\rightarrow\infty}\frac{N(\lambda)-c_{1}\lambda^{d}}{\lambda^{d-1}}=c_{2}=\lim_{\lambda\rightarrow\infty}\frac{M(\lambda)-c_{1}\lambda^{d}}{\lambda^{d-1}},$
where $c_{1},c_{2}$ are non-zero real numbers.
Proposition 7.5.
Let $c_{1}$ and $c_{2}$ be non-zero real numbers and $k\in\mathbb{Z}^{\geq 1}$. Define $N,M:\mathbb{R}\longrightarrow\mathbb{R}$ as
$N(\lambda)$ $:=c_{1}[\lambda]^{k}+c_{2}[\lambda]^{k-1}+\alpha([\lambda])\qquad\textnormal{and}\qquad M(\lambda):=c_{1}\lambda^{k}+c_{2}\lambda^{k-1}+\beta(\lambda),$
where $\alpha$ and $\beta$ are polynomials of degree at most $k-2$ if $k\geq 2$ and they are zero if $k=1$. Then $N\sim_{2}M$.
Proof.
This follows by unwinding the definitions and using the fact $\lim_{\lambda\longrightarrow\infty}\frac{[\lambda]}{\lambda}=1.$
Proposition 7.6.
Let $\widetilde{M}:\mathbb{R}\longrightarrow\mathbb{R}$ be defined as
$\widetilde{M}(\gamma):=c_{1}\gamma^{d}+c_{2}\gamma^{d-1},$
where $d\in\mathbb{Z}^{\geq 1}$ and $c_{1},c_{2}$ are non-zero real numbers. Let $\widetilde{N}:\mathbb{R}\longrightarrow\mathbb{R}$ be a function such that $\widetilde{N}\sim_{2}\widetilde{M}$. Let us define $\gamma:\mathbb{R}\longrightarrow\mathbb{R}$ as
$\gamma(\lambda):=a\lambda+b,$
where $a\in\mathbb{R}^{>0}$ and $b\in\mathbb{R}$. Define $N,M:\mathbb{R}\longrightarrow\mathbb{R}$ as
$N(\lambda):=\widetilde{N}(\gamma(\lambda))\qquad\textnormal{and}\qquad M(\lambda):=c_{1}a^{d}\lambda^{d}+(bdc_{1}+c_{2})a^{d-1}\lambda^{d-1}.$
If $(bdc_{1}+c_{2})\neq 0$, then $N\sim_{2}M$.
Proof.
The proof will follow by unwinding the definition of a limit. First of all, we note that
$\frac{N(\lambda)}{\lambda^{d}}$ $=\Big{(}\frac{\widetilde{N}(\gamma(\lambda))}{\gamma(\lambda)^{d}}\Big{)}\Big{(}\frac{\gamma(\lambda)^{d}}{\lambda^{d}}\Big{)}.$ (57)
We now note that
$\lim_{\lambda\longrightarrow\infty}\Big{(}\frac{\widetilde{N}(\gamma(\lambda))}{\gamma(\lambda)^{d}}\Big{)}$ $=c_{1}\qquad\textnormal{and}$ (58)
$\lim_{\lambda\longrightarrow\infty}\Big{(}\frac{\gamma(\lambda)^{d}}{\lambda^{d}}\Big{)}$ $=a^{d}.$ (59)
Equation (58) follows from the fact that $\lim_{\lambda\longrightarrow\infty}\frac{\widetilde{N}(\lambda)}{\lambda^{d}}=c_{1}$ and the fact that $\gamma(\lambda)$ is a monotonically increasing function of $\lambda$ and applying the definition of a limit. Equations (58), (59) and (57) imply that
$\lim_{\lambda\longrightarrow\infty}\frac{N(\lambda)}{\lambda^{d}}=c_{1}a^{d}.$ (60)
Next, we note that
$\frac{N(\lambda)-c_{1}a\lambda^{d}}{\lambda^{d-1}}$ $=\Big{(}\frac{\widetilde{N}(\gamma(\lambda))-c_{1}\gamma(\lambda)^{d}+c_{1}\gamma(\lambda)^{d}-c_{1}a^{d}\lambda^{d}}{\gamma(\lambda)^{d-1}}\Big{)}\Big{(}\frac{\gamma(\lambda)^{d-1}}{\lambda^{d-1}}\Big{)}$
$=\Big{(}\frac{\widetilde{N}(\gamma(\lambda))-c_{1}\gamma(\lambda)^{d}}{\gamma(\lambda)^{d-1}}\Big{)}\Big{(}\frac{\gamma(\lambda)^{d-1}}{\lambda^{d-1}}\Big{)}+\Big{(}\frac{c_{1}\gamma(\lambda)^{d}-c_{1}a^{d}\lambda^{d}}{\lambda^{d-1}}\Big{)}$
$\implies\lim_{\lambda\longrightarrow\infty}\Big{(}\frac{N(\lambda)-c_{1}a\lambda^{d}}{\lambda^{d-1}}\Big{)}$ $=\lim_{\lambda\longrightarrow\infty}\Big{(}\frac{\widetilde{N}(\gamma(\lambda))-c_{1}\gamma(\lambda)^{d}}{\gamma(\lambda)^{d-1}}\Big{)}a^{d-1}+\lim_{\lambda\longrightarrow\infty}\Big{(}\frac{c_{1}\gamma(\lambda)^{d}-c_{1}a^{d}\lambda^{d}}{\lambda^{d-1}}\Big{)}$
$=a^{d-1}c_{2}+a^{d-1}bdc_{1}.$ (61)
The first limit on the right hand side of equation (61) is computed using the fact that $\gamma(\lambda)$ is a monotonically increasing function of $\lambda$ and the fact that $\widetilde{N}\sim_{2}\widetilde{N}$ and then applying definition of a limit. The second limit can be computed immediately by considering the binomial expansion of $(a\lambda+b)^{d}$. Equations (60) and (61) imply that $N\sim_{2}M$, whence the claim is proved. ∎
We are now ready to prove the main claim.

### v2

@@ 1. Introduction
$N_{M}(\lambda)$ $:=\#\{\textnormal{eigenvalues of $\Delta_{M}$}\leq\lambda^{2}\}.$
@@ 3. Setup and notation
$\Lambda_{M}$ $:=\{\sqrt{\alpha_{M}(k)}:k=0,1,2,\ldots\},\qquad\textnormal{where}\qquad\alpha_{M}(k):=A_{M}k^{2}+B_{M}k+C_{M}.$ (5)
@@
Assume that the multiplicity of these eigenvalues are given by a degree d1d-1 polynomial $\mathsf{R}_{M}(k)$. Define
@@
where $\varphi(t)$ is a polynomial of degree at most d3d-3 and $C$ is some constant.
@@ 4. A lattice counting result by Iosevich and Wyman
Using this lattice counting theorem, we are going to prove the following result.
@@
Let $d_{1},d_{2},\ldots,d_{n}\in\mathbb{Z}^{\geq 2}$ and $n\geq 2$. Let $y_{1},\ldots,y_{n}$ be fixed $n$ non negative real numbers and $P_{1},P_{2},\ldots P_{n}$ be fixed $n$ positive real numbers. Then
@@
We first prove an auxiliary result that combined with Theorem 4.1 will immediately imply Theorem 4.2. First we make the following notation: define the norm $p$ on $\mathbb{R}^{n}$ by
@@
The crucial fact that will eventually be used (in order to apply Theorem 4.1) is that the ellipsoid $p^{-1}(1)$ has non vanishing Gaussian curvature. We will denote $B_{p}$ to be the unit ball in $\mathbb{R}^{n}$ with respect to the norm $p$, i.e.
$B_{p}$ $:=\{\xi\in\mathbb{R}^{n}:p(\xi)^{2}\leq 1\}.$
Lemma 4.3.
With notation as above, define
$N(\lambda)$ $:=\sum_{{m\in(\mathbb{Z}_{\geq 0}^{n}+y)}\atop{p(m)^{2}\leq\lambda^{2}}}m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}\qquad\textnormal{and}\qquad N_{2}(\lambda):=\sum_{{m\in(\mathbb{Z}^{n}+y)\cap\mathbb{R}_{\geq 0}^{n}}\atop{p(m)^{2}\leq\lambda^{2}}}m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}.$
Then,
$N_{2}(\lambda)$ $=N(\lambda)+O(\lambda^{d-2}).$ (10)
Proof: Note that we can write $N(\lambda)$ as
$N(\lambda)$ $=\sum_{m\in\mathbb{Z}^{n}_{\geq 0}+y}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1},$
where $\chi$ denotes the characteristic function. Let us define the following sets for $j\in\{1,2,\ldots,n\}$
$S$ $:=\big(\mathbb{Z}^{n}+y\big)\cap\mathbb{R}_{\geq 0}^{n},\qquad H_{j}:=\{x\in\mathbb{R}^{n}_{\geq 0}:x_{j}\in[0,y_{j}+2)\},$
$L_{j}$ $:=\big(\mathbb{Z}^{n}+y\big)\cap H_{j}\qquad\textnormal{and}\qquad A_{j}:=\{m\in\mathbb{Z}_{\geq 0}^{n}+y:m_{j}-y_{j}\in\{0,1\}\}.$
Using the fact that $y_{j}\geq 0$ for all $j$, the following two set theoretic identities can easily be verified
$S$ $=\Big(\mathbb{Z}^{n}_{\geq 2}+y\Big)\bigsqcup\bigcup_{j=1}^{n}L_{j}\qquad\textnormal{and}$ (11)
$\big(\mathbb{Z}_{\geq 0}^{n}+y\big)$ $=\big(\mathbb{Z}_{\geq 2}^{n}+y\big)\bigsqcup\bigcup_{j=1}^{n}A_{j},$ (12)
where $\sqcup$ denotes disjoint union. Note that $y_{j}\geq 0$ is used only to prove (11) (for (12) that fact is not relevant). Let us now rewrite our given functions using the notation we have introduced. We note that
$N(\lambda)$ $=\sum_{m\in\mathbb{Z}^{n}_{\geq 0}+y}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}\qquad\textnormal{and}\qquad$ (13)
$N_{1}(\lambda)$ $=\sum_{m\in\mathbb{Z}^{n}_{\geq 2}+y}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}$ (14)
Note that since $y_{i}\geq 0$ for all $i\in\{1,2,\ldots,n\}$, we conclude that the summand on the right hand side of equation (13) is non negative for all $m$ in the domain of the summation. Furthermore, $N_{1}$ is defined by summing up the same function over a smaller set. Hence, using (12), we conclude that
$N_{1}(\lambda)\leq N(\lambda)$ $\leq N_{1}(\lambda)+\sum_{j=1}^{n}\sum_{m\in A_{j}}\chi_{B_{p}}(m/\lambda)m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}.$ (15)
Note that $y_{i}\geq 0$ is crucial for this argument since that ensures that the summand is non negative. Next, let us define for all $j\in\{1,\ldots,n\}$ and for all $\sigma\in\{0,1\}$
$\mathsf{R}_{j\sigma}$ $:=\sum_{{m\in\mathbb{Z}^{n}_{\geq 0}+y}\atop{m_{j}=y_{j}+\sigma}}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}m_{2}^{d_{2}-1}\ldots m_{n}^{d_{n}-1}.$ (16)
We now note that the second term on the right hand side of (15) is precisely equal to
$\sum_{j=1}^{n}\sum_{\sigma=0}^{1}\mathsf{R}_{j\sigma}$
Hence, equation (15) can be rewritten as
$0\leq N(\lambda)-N_{1}(\lambda)$ $\leq\sum_{j=1}^{n}\sum_{\sigma=0}^{1}\mathsf{R}_{j\sigma}$
$\implies\big|N(\lambda)-N_{1}(\lambda)\big|$ $\leq\sum_{j=1}^{n}\sum_{\sigma=0}^{1}\mathsf{R}_{j\sigma}.$ (17)
Next, let us study the function $N_{2}$. We note that
$N_{2}(\lambda)$ $=\sum_{m\in S}\chi_{B_{p}}(m/\lambda)m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}.$ (18)
Using equation (11) and the non-negativity of the summand in the definition of $N_{2}$, we conclude that
$0\leq N_{2}(\lambda)-N_{1}(\lambda)$ $\leq\sum_{j=1}^{n}\sum_{m\in L_{j}}\chi_{B_{p}}(m/\lambda)m_{1}^{d_{1}-1}m_{2}^{d_{2}-1}\ldots m_{n}^{d_{n}-1}$
$\implies\big|N_{2}(\lambda)-N_{1}(\lambda)\big|$ $\leq\sum_{j=1}^{n}\sum_{m\in L_{j}}\chi_{B_{p}}(m/\lambda)m_{1}^{d_{1}-1}m_{2}^{d_{2}-1}\ldots m_{n}^{d_{n}-1}.$ (19)
Hence, to prove the Lemma (namely equation (10)) it suffices to show that it suffices to show that
$\mathsf{R}_{j\sigma}=O(\lambda^{d-2})\qquad\forall~~j\in\{1,\ldots,n\}\qquad\textnormal{and}\qquad\forall\sigma\in\{0,1\}\qquad\textnormal{and}$ (20)
$\sum_{m\in L_{j}}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}=O(\lambda^{d-2})\qquad\forall~~j\in\{1,\ldots,n\}.$ (21)
To see why that is so, we note that if we do manage to prove the above two claims, then by (17), we will conclude that $N(\lambda)-N_{1}(\lambda)\in O(\lambda^{d-2})$ while using (17), we will conclude that $N_{2}(\lambda)-N_{1}(\lambda)\in O(\lambda^{d-2})$. Using the fact that if $f,g\in O(\lambda^{k})$ then fgf-g also belongs to $O(\lambda^{k})$, we conclude that $N_{2}-N=(N_{2}-N_{1})-(N-N_{1})$ is in $O(\lambda^{d-2})$. Let us now prove equations (20) and (21).
      We start by proving (20). Without loss of generality, it is enough to prove equation (20) for $j=1$. We show that now. In the subsequent calculations, we will make a standard abuse of notation often used by analysts: if the letter $C$ appears in two different lines, then they usually refer to two different constants. Keeping this in mind, we note that
$\mathsf{R}_{1\sigma}$ $=(y_{1}+\sigma)^{d_{1}-1}\sum_{{m_{2}-y_{2},\ldots,m_{n}-y_{n}\in\mathbb{Z}_{\geq 0}}\atop{m_{1}=y_{1}+\sigma}}\chi_{B_{p}}(m/\lambda)\ m_{2}^{d_{2}-1}\ldots m_{n}^{d_{n}-1},$ (22)
$\implies\mathsf{R}_{1\sigma}$ $\leq C\sum_{{m_{2}-y_{2},\ldots,m_{n}-y_{n}\in\mathbb{Z}_{\geq 0}}\atop{m_{1}=y_{1}+\sigma}}\chi_{B_{p}}(m/\lambda)\ p(m)^{d-d_{1}-n+1}.$ (23)
Let us justify (23). Note that for any $j\in\{2,\ldots,n\}$, we have
$m_{j}^{2(d_{j}-1)}$ $=\Big(\Big(\frac{1}{P_{j}}\Big)^{2(d_{j}-1)}\Big)\Big((P_{j}m_{j})^{2}\Big)^{d_{j}-1}.$
Hence,
$m_{j}^{2(d_{j}-1)}$ $\leq\Big(\frac{1}{P_{j}}\Big)^{2(d_{j}-1)}\Big(P_{1}^{2}m_{1}^{2}+\ldots+P_{n}m_{n}^{2}\Big)^{d_{j}-1}\qquad\forall~~j\in\{2,\ldots,n\}.$ (24)
Multiplying the inequalities (24) from $j=2$ to $n$ gives us
$m_{2}^{2(d_{2}-1)}m_{3}^{2(d_{3}-1)}\ldots m_{n}^{2(d_{n}-1)}$ $\leq C\Big(P_{1}^{2}m_{1}^{2}+\ldots+P_{n}m_{n}^{2}\Big)^{d-d_{1}-n+1}.$ (25)
Equation (22) combined with equation (25) gives us equation (23) (we remind the reader that the $C$ in equations (25) and (23) refer to two different constants). Next, we will use the following standard fact about the asymptotic behaviour about the number of lattice points inside a ball of radius $\lambda$. Define
$M(\lambda)$ $:=\sum_{{m_{1}-y_{1},m_{2}-y_{2},\ldots,m_{n}-y_{n}\in\mathbb{Z}}\atop{P_{1}m_{1}^{2}+\ldots+P_{n}m_{n}^{2}\leq\lambda^{2}}}\mathbf{1}.$ (26)
Then
$M(\lambda)$ $\leq C\lambda^{n}.$ (27)
Although this is a standard fact, we have included a proof of equation (27) in section 8.
Using equation (27), we conclude that
$\mathsf{R}_{1\sigma}$ $\leq C\lambda^{d-d_{1}-n+1}\sum_{{m_{2}-y_{2},\ldots,m_{n}-y_{n}\in\mathbb{Z}_{\geq 0}}\atop{m_{1}=y_{1}+\sigma}}\chi_{B_{p}}(m/\lambda)$
$=C\lambda^{d-d_{1}-n+1}\sum_{\begin{subarray}{c}m_{2}-y_{2},m_{3}-y_{3},\ldots,m_{n}-y_{n}\in\mathbb{Z}\\ P_{2}m_{2}^{2}+\ldots+P_{n}m_{n}^{2}\leq\lambda^{2}-P_{1}m_{1}^{2}\\ m_{1}=y_{1}+\sigma\end{subarray}}\mathbf{1}$
$\leq C\lambda^{d-d_{1}-n+1}\sum_{\begin{subarray}{c}m_{2}-y_{2},m_{3}-y_{3},\ldots,m_{n}-y_{n}\in\mathbb{Z}\\ P_{2}m_{2}^{2}+\ldots+P_{n}m_{n}^{2}\leq\lambda^{2}\end{subarray}}\mathbf{1}$
$\leq C\lambda^{d-d_{1}-n+1}\lambda^{n-1}$ (28)
$=C\lambda^{d-d_{1}}$
$=O(\lambda^{d-2})\qquad\textnormal{(since $d_{1}\geq 2$).}$ (29)
Equation (28) follows from equation (27) (with $n$ replaced by n1n-1). Equation (29) follows from the fact that $d_{1}\geq 2$. This proves (20). It now remains to show (21), which we now prove. Again, it suffices to prove (21) for $j=1$. Note that
$\sum_{m\in L_{1}}\chi_{B_{p}}(m/\lambda)\ m_{1}^{d_{1}-1}\ldots m_{n}^{d_{n}-1}$ $\leq(y_{1}+2)^{d_{1}-1}\sum_{m\in L_{1}}\chi_{B_{p}}(m/\lambda)m_{2}^{d_{2}-1}\ldots m_{n}^{d_{n}-1}$
$\leq C\ \lambda^{d-d_{1}-n+1}\sum_{m\in L_{1}}\chi_{B_{p}}(m/\lambda)$ (30)
$=C\ \lambda^{d-d_{1}-n+1}\lambda^{n-1}$ (31)
$=C\ \lambda^{d-d_{1}}$
$=O(\lambda^{d-2}).$ (32)
Equation (30) follows from equation (25). Equation (31) follows from (27) (with $n$ replaced with n1n-1). Finally, equation (32) follows from the fact that $d_{1}\geq 2$. This proves Lemma 4.3. ∎
Proof of Theorem 4.2: From Lemma 4.3:
$N_{2}(\lambda)=N(\lambda)+O(\lambda^{d-2})$ (33)
and Theorem 4.1 gives:
$N_{2}(\lambda)=C\lambda^{d}+O(\lambda^{d-1-\frac{n-1}{n+1}}).$ (34)
Equations (33) and (34) combine to give: $N(\lambda)=C\lambda^{d}+O(\lambda^{d-1-\frac{n-1}{n+1}})$ and that is precisely the statement of Theorem 4.2. ∎
@@ 5. Error can be polynomially improved for products of $\mathsf{W}$-manifolds: proof
Note that the sum total of the powers of all the $m_{i}$ is dnd-n.
@@
where the sum of the total power $e_{1}+\ldots e_{n}$ is at most dn2d-n-2 (due to the condition 35). In particular, the sum total of the powers of $m_{i}$ in the summand of each of the terms of $T_{2}$ is at least $2$ less than the sum total of the powers of the $m_{i}$ in the summand of $T_{1}$.
@@
Similarly, we again apply theorem 4.2 and the binomial theorem to each of the terms of $T_{2}$. Doing that, we conclude that $T_{2}$ will be given by the right hand side of (42) with $d$ replaced by d2d-2, i.e.
@@ 6.1. Odd dimensional spheres
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree 2d22d-2. Now,
@@ 6.2. Even dimensional spheres
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree 2d12d-1. Now,
@@ 6.3. Odd dimensional real projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree 2d22d-2. Now,
@@ 6.4. Even dimensional real projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree 2d12d-1. Now,
@@ 6.5. Complex projective spaces
Note that $\mathsf{R}_{M}(k)$ is a polynomial of degree 2d12d-1. Now,
@@ 6.6. Quaternionic projective spaces
Let us now move on to the quaternionic projective spaces $M=\mathbb{HP}^{d-1}$, with $d\geq 2$. We are referring to the (quaternionic) dimension as d1d-1 as opposed to $d$ in order to make it easy to follow the notation of [4]. Notice that the real dimension of $\mathbb{HP}^{d-1}$ is 4d44d-4. Following [4, page 10, theorem 5.2], we conclude that
@@
Notice that $\mathsf{R}_{M}(k)$ is a polynomial of degree 4d54d-5. Now,
@@ 7. Error is sharp for a single $\mathsf{W}$-manifold: proof
where $\mathcal{R}_{k-2}(n)$ is a polynomial in $n$ of degree at most k2k-2.
@@
$2^{k}-1^{k}=k\cdot 1^{k-1}+\frac{k(k-1)}{2}\cdot 1^{k-2}+\cdots+1.$
@@
Next, we prove the following result.
Lemma 7.2.
Let $a\in\mathbb{R}_{>0}$, $b\in\mathbb{R}$ and $y\in\mathbb{R}_{\geq 0}$. For $\lambda$ sufficiently large and positive, define
$M(\lambda)$ $=\frac{[\sqrt{a\lambda^{2}+b}-y]^{d}}{d}+(y+\frac{1}{2})[\sqrt{a\lambda^{2}+b}-y]^{d-1}.$
Then
$\lim_{\lambda\longrightarrow+\infty}\frac{M(\lambda)}{\lambda^{d}}$ $=\frac{a^{\frac{d}{2}}}{d}.$ (81)
Furthermore, the error term $\mathcal{E}(\lambda)$ can not be $O(\lambda^{d-1-\varepsilon})$ for any positive $\varepsilon$, where
$\mathcal{E}(\lambda)$ $:=M(\lambda)-\frac{a^{\frac{d}{2}}}{d}\lambda^{d}.$
Proof: Equation (81) is immediate using the fact that
$A+B-1$ $\leq[A+B]\leq A+B+1.$ (82)
An application of (82), binomial theorem and the squeeze rule immediately gives us (81). We now prove the second assertion. We will prove by contradiction. Suppose $\mathcal{E}(\lambda)$ was in $O(\lambda^{d-1-\varepsilon})$ for some $\varepsilon>0$. Let us define the following discrete subset of $\mathbb{R}$ given by
$\Lambda$ $:=\Big\{\lambda\in\mathbb{R}_{>0}:a\lambda^{2}+b>0,\sqrt{a\lambda^{2}+b}-y\in\mathbb{Z}_{\geq 0}\Big\}.$
Define $\gamma:\Lambda\longrightarrow\mathbb{Z}_{>0}$, given by
$\gamma(\lambda)$ $:=\sqrt{a\lambda^{2}+b}-y.$ (83)
To lighten the notation, we will abbreviate $\gamma(\lambda)$ by $\gamma$; however, it should always be understood that $\gamma$ depends on $\lambda$ via equation (83). Hence, if the error term was in $O(\lambda^{d-1-\varepsilon})$, then it would imply that there is some constant $C>0$ such that for all $\lambda\in\Lambda$, sufficiently large, we have
$\Big|\Big(y+\frac{1}{2}\Big)\gamma^{d-1}+\Big(\frac{\gamma^{d}}{d}-\frac{a^{\frac{d}{2}}}{d}\lambda^{d}\Big)\Big|$ $\leq C\lambda^{d-1-\varepsilon}.$ (84)
Using the fact that $|A|-|B|\leq|A+B|$ (note that $|A+B-B|\leq|A+B|+|B|$ which implies what we have written), we conclude that
$\Big|\Big(y+\frac{1}{2}\Big)\gamma^{d-1}\Big|-\Big|\frac{\gamma^{d}}{d}-\frac{a^{\frac{d}{2}}}{d}\lambda^{d}\Big|$ $\leq C\lambda^{d-1-\varepsilon}.$ (85)
Here we set
$A:=\Big(y+\frac{1}{2}\Big)\gamma^{d-1}\qquad\textnormal{and}\qquad B:=\Big(\frac{\gamma^{d}}{d}-\frac{a^{\frac{d}{2}}}{d}\lambda^{d}\Big).$
Now we divide both sides of equation (85) and take the limit as $\lambda$ tends to infinity (and $\lambda\in\Lambda$). The right hand side will obviously be zero. We first note that
$\lim_{\begin{subarray}{c}\lambda\longrightarrow\infty\\ \lambda\in\Lambda\end{subarray}}\frac{\gamma^{d}}{d}-\frac{a^{\frac{d}{2}}}{d}\lambda^{d}$ $=0.$
This follows by taking a binomial expansion of $\gamma$ and noticing that after subtracting the leading term (namely $\frac{a^{\frac{d}{2}}}{d}\lambda^{d}$), we are left with a terms of order $\lambda^{d-2}$ or less. Hence, what remains is the first term on the left hand side of equation (85). We note that
$\lim_{\begin{subarray}{c}\lambda\longrightarrow\infty\\ \lambda\in\Lambda\end{subarray}}\Big(y+\frac{1}{2}\Big)\frac{\gamma^{d-1}}{\lambda^{d-1}}$ $=\Big(y+\frac{1}{2}\Big).$
Since $y$ is non negative, the above limit is positive. This gives us a contradiction. ∎